# A Unified Framework for Spatial and Temporal Treatment Effect Boundaries: Theory and Identification

### v1

@@ 7.1 Main Contributions
Theoretical unification: We formalize spatial and temporal treatment effect boundaries as structural parameters arising from a common diffusion process. Under the proposed model, boundaries satisfy $d^{*}/\tau^{*}\approx\lambda^{2}/\sqrt{\delta}$, linking spatial reach to temporal persistence through decay parameters.
@@ 7.5 Concluding Remarks
As quasi-experimental methods continue to advance, incorporating spatial and temporal dynamics explicitly—rather than treating them as nuisances—will become increasingly important. Our framework provides tools for this next generation of empirical work, where understanding not just whether policies work, but where, when, and for how long they operate, is central to informed decision-making.
The boundary is not where analysis ends—it is where understanding begins.

### v2

@@ 8.1 Main Contributions
Theoretical unification: We formalize spatial and temporal treatment effect boundaries as structural parameters arising from a common diffusion process. Under the proposed model, boundaries satisfy $d^{*}/\tau^{*}=\lambda\sqrt{\delta}\cdot c$ where $c$ depends on detection thresholds (typically $c\approx 3.32$ for standard 10%/50% thresholds), linking spatial reach to temporal persistence through decay parameters $(\delta,\lambda)$.
@@ 8.3.1 Functional Form Assumptions
As demonstrated in Kikuchi (2024), diffusion-based approaches provide theoretical guidance for spatial boundary detection. The current paper extends this framework by incorporating temporal dynamics and deriving over-identification tests linking spatial reach to temporal persistence.
@@ 8.5 Concluding Remarks
As quasi-experimental methods continue to advance, incorporating spatial and temporal dynamics explicitly – rather than treating them as nuisances – will become increasingly important. Our framework provides tools for this next generation of empirical work, where understanding not just whether policies work, but where, when, and for how long they operate, is central to informed decision-making.
— The boundary is not where analysis ends – it is where understanding begins.